# Open-vocabulary Temporal Action Localization using VLMs

### v2

@@ I Introduction
In this paper, we introduce a pipeline for Open-vocabulary video action localization (OVAL), which does not require pre-training and can address action labels in an open-vocabulary form (Fig. 1). Our approach is inspired by a recently proposed method called “Prompting with Iterative Visual Optimization (PIVOT)”[8]. The original PIVOT aims to utilize Vision-Language Models (VLM) to reason an embodied action by framing the problem as iterative visual question answering. In this paper, we propose a simple yet effective approach to apply PIVOT to find a temporal segment of actions. Specifically, we sample video frames into a concatenated image with frame index labels, making a VLM guess the frame closest to the start and end of an action. Iterating this process by narrowing the sampling time window results in finding specific frames when the action started and ended, thus achieving action localization. Although the recent advancements in VLM technology has supported the development of OVAL [9, 10, 11, 12, 13, 14, 15], to the best of our knowledge, there are no existing methods that utilize off-the-shelf VLMs directly.
@@
II Method
II-A Core Pipeline
@@ III-B Task Definition
III-C T-PIVOT for Recognizing Seamless Action Transitions
By applying the T-PIVOT method described in Section II, we estimated the timing of task transitions within the video as follows. First, the start time of each task, $T_{i\_start}$ ($i=1\ldots N$), is estimated using the T-PIVOT in parallel processing. The end time of each task, $T_{i\_end}$ ($i=1\ldots N$), is similarly estimated. This approach yields an estimated series of start times and end times for each task. In this experiment, we assume that tasks are continuously present in the video. Therefore, the end time of one task is identical to the start time of the subsequent task. To ensure this, we calculate this time as the midpoint between the estimated end time of the current task and the estimated start time of the next task:
@@ III-C T-PIVOT for Recognizing Seamless Action Transitions
III-D Action-order Aware Prompting
@@
III-E Performance Across Different Prompting Styles
@@
III-F Comparison with Existing Methods

### v3

@@ I Introduction
In this paper, we introduce a pipeline for Open-vocabulary video action localization (OVAL), which does not require pre-training and can address action labels in an open-vocabulary form (Fig. 1). Our approach is inspired by a recently proposed method called “Prompting with Iterative Visual Optimization (PIVOT)”[8]. The original PIVOT aims to utilize Vision-Language Models (VLM) to reason an embodied action by framing the problem as iterative visual question answering. In this paper, we propose a simple yet effective approach to apply PIVOT to find a temporal segment of actions. Specifically, we sample video frames into a concatenated image with frame index labels, making a VLM guess the frame closest to the start and end of an action. Iterating this process by narrowing the sampling time window results in finding specific frames when the action started and ended, thus achieving action localization. Although the recent advancements in VLM technology has supported the development of OVAL [9, 10, 11, 12, 13, 14], to the best of our knowledge, there are no existing methods that utilize off-the-shelf VLMs directly.
@@
II Related Work
II-A Model-based Action Localization
Model-based temporal action localization can be broadly categorized into anchor-based methods and boundary-based methods. Anchor-based methods, also known as top-down methods, mainly use multi-scale temporal sliding windows to detect pre-trained actions within untrimmed videos [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. In contrast, boundary-based methods, also known as bottom-up methods, initially detect the temporal boundaries of actions, followed by further analysis of the proposal regions estimated from these boundaries [31, 6, 7, 27, 32, 33]. These methods typically involve a two-stage pipeline consisting of temporal proposal generation and action classification [25, 27, 28, 26], while end-to-end methods have also been proposed [34, 19, 35, 36]. The proposed method follows a boundary-based approach, as it aims to detect the start and end boundaries of actions. Additionally, unlike other methods, our VLM-based method accommodates open-vocabulary queries. This allows for the addition of new actions without the need for data re-collection or re-training.
II-B Open-vocabulary Action Localization
Open-vocabulary action localization can be broadly categorized into two approaches. The first approach outputs the closest matching action label from a pre-trained action set, while the latter is class-label-free, aiming to localize actions specified by a free-text query. In the former approach, LLMs or high-level scene understanding are used to bridge the gap between the query and pre-trained action labels [10, 9]. In a typical implementation of the latter approach, actions are first detected and then the corresponding action segment is determined based on the relationship between video features and the query [14]. Pre-trained multimodal models are often used to achieve these objectives [12, 13, 37, 38, 11]. The proposed method follows the class-label-free approach, leveraging off-the-shelf models without additional training. Its unique advantage lies in being training-free and using VLMs to directly match action detection with language in a single stage.
III Method
III-A Core Pipeline
@@ IV-B Task Definition
IV-C T-PIVOT for Recognizing Seamless Action Transitions
By applying the T-PIVOT method described in Section III, we estimated the timing of task transitions within the video as follows. First, the start time of each task, $T_{i\_start}$ ($i=1\ldots N$), is estimated using the T-PIVOT in parallel processing. The end time of each task, $T_{i\_end}$ ($i=1\ldots N$), is similarly estimated. This approach yields an estimated series of start times and end times for each task. In this experiment, we assume that tasks are continuously present in the video. Therefore, the end time of one task is identical to the start time of the subsequent task. To ensure this, we calculate this time as the midpoint between the estimated end time of the current task and the estimated start time of the next task:
@@ IV-C T-PIVOT for Recognizing Seamless Action Transitions
IV-D Action-order Aware Prompting
@@
IV-E Performance Across Different Prompting Styles
@@
IV-F Comparison with Existing Methods